# Spatial Distribution-Shift Aware Knowledge-Guided Machine Learning

### v1

@@ Abstract
Given inputs of diverse soil characteristics, and climate data gathered from various regions, we aimed to build a model to predict accurate land emissions. The problem is important since accurate quantification of the carbon cycle in agroecosystems is crucial for mitigating climate change and ensuring sustainable food production. Predicting accurate land emissions is challenging due to since calibrating heterogeneous nature of soil properties, moisture, and environmental conditions is hard at decision-relevant scales. Traditional approaches do not adequately estimate land emissions due to location-independent parameters failing to leverage the spatial heterogeneity and also require large datasets. To overcome these limitations, we proposed Spatial Distribution-Shift Aware Knowledge-Guided Machine Learning (SDSA-KGML) which leverage location-dependent parameters which accounts significant spatial heterogeneity in soil moisture from multiple sites within the same region. Experimental results demonstrate that SDSA-KGML models achieve higher local accuracy for the specified states in the Midwest Region.
@@ 3 Spatial Variability-Aware Proposed Approach
Future Work: We plan to add novel partitioning techniques and new datasets (Kumar et al. 2015) for detailed experiments to better capture spatial variability (Ghosh, Sharma et al. 2024b, a; Ghosh et al. 2024a, b, 2023, 2022, 2017; Yang et al. 2025). In addition, we also plan to incorporate Task-Adaptive Meta Learning (Liu et al. 2023) for model generalization corresponding to any new tasks.
@@ Acknowledgments
This material is based on work supported by the USDA under Grant No. 2023-67021-39829, the National Science Foundation under Grant No. 1901099, and the USDOE Office of Energy Efficiency and Renewable Energy under FOA No. DE-FOA0002044. We also thank Kim Kofolt and the Spatial Computing Research Group for their valuable comments and contributions..

### v2

@@ Abstract
Given inputs of diverse soil characteristics and climate data gathered from various regions, we aimed to build a model to predict accurate land emissions. The problem is important since accurate quantification of the carbon cycle in agroecosystems is crucial for mitigating climate change and ensuring sustainable food production. Predicting accurate land emissions is challenging since calibrating the heterogeneous nature of soil properties, moisture, and environmental conditions is hard at decision-relevant scales. Traditional approaches do not adequately estimate land emissions due to location-independent parameters failing to leverage the spatial heterogeneity and also require large datasets. To overcome these limitations, we proposed Spatial Distribution-Shift Aware Knowledge-Guided Machine Learning (SDSA-KGML), which leverages location-dependent parameters that account for significant spatial heterogeneity in soil moisture from multiple sites within the same region. Experimental results demonstrate that SDSA-KGML models achieve higher local accuracy for the specified states in the Midwest Region.
@@ 3 Spatial Variability-Aware Proposed Approach
Future Work: We plan to add novel partitioning techniques and new datasets (Kumar et al. 2015) for detailed experiments to better capture spatial variability (Ghosh, Sharma et al. 2024b, a; Ghosh et al. 2024a, b, 2023, 2022; ghosh2017video; Yang et al. 2025). In addition, we also plan to incorporate Task-Adaptive Meta Learning (Liu et al. 2023) for model generalization corresponding to any new tasks.
@@ Acknowledgments
This material is based on work supported by the USDA under Grant No. 2023-67021-39829, the National Science Foundation under Grant No. 1901099, the USDOE Office of Energy Efficiency and Renewable Energy under FOA No. DE-FOA0002044, and USDA under Grant No. 2021-51181-35861. We also thank Kim Kofolt and the Spatial Computing Research Group for their valuable comments and contributions.